A 2010 study asserts that the number of hours that the average college student studies each week has been steadily dropping (The Boston Globe, July 4, 2010). In fact, the researchers state that, in the U.S., today’s undergraduates study an average of 14 hours per week. Suppose an administrator at a local university wants to show that the average study time of students at his university differs from the national average. He takes a random sample of 35 students at is university and finds that the average number of hours spent studying per week is 16.3. Assume that the population standard deviation is 7.2 hours. At 0.05 level of significance, what is the test statistics for testing the hypotheses H0: µ = 14 versus H1: µ ≠ 14?

Select one:

2.05

1.89

2.22

1.96

Question 6

Not yet answered

Points out of 1.00

Flag question

Question text

Suppose the Z test statistic in a one-tailed test is 1.76. What is the p-value?

Select one:

0.0392

0.0212

0.0424

0.0588

none of these are correct
Answer #1

The first test is two tailed with population sd known.

The test is z test.

Test Statistic :

z=(​​​​​​

obsd z=(16.3-14)/(7.2/​​​​​​)

=1.89

2nd option is correct.

QUESTION 6

Cinsider the test is right tailed,

p value=P(Z>1.76|Z follow Normal(0,1))

=1-P(Z1. 76)=1-0.9607961=0.0392

First option is correct.